### v2

# Distribution-free data-driven smooth tests without $\chi^{2}$

Abstract
This article demonstrates how recent developments in the theory of empirical processes allow us to construct a new family of asymptotically distribution-free smooth test statistics. Their distribution-free property is preserved even when the parameters are estimated, model selection is performed, and the sample size is only moderately large. A computationally efficient alternative to the classical parametric bootstrap is also discussed.
Keywords: goodness-of-fit tests, smooth tests, distribution-freeness
1 Introduction
Let $X$ be a continuous random variable with cumulative distribution function (CDF) $Q$ and probability density function (PDF) $q$ with support $\mathcal{X}\subseteq\mathbb{R}$. We are interested in assessing whether the unknown distribution $Q$ belongs to a family of distribution functions $G_{\bm{\beta}}$, with PDF $g_{\bm{\beta}}$, where $\bm{\beta}\in\mathcal{B}\subseteq\mathbb{R}^{p}$. To tackle this problem, we may consider an alternative model of the form
$g_{\bm{\beta}}(x)\Big{\{}1+\sum_{j=1}^{m}\theta_{j\bm{\beta}}h_{j\bm{\beta}}(x)\Big{\}}$ (1)
which incorporates the null density, $g_{\bm{\beta}}$, as a special case. The functions $\{h_{j\bm{\beta}}\}_{j=1}^{\infty}$ form an orthonormal basis in $L^{2}(G_{\bm{\beta}})$; hence
$\int_{\mathcal{X}}h_{i\bm{\beta}}(x)h_{j\bm{\beta}}(x)dG_{\bm{\beta}}(x)=\mathds{1}_{\{i=j\}},$
where $\mathds{1}_{\{\cdot\}}$ denotes the indicator function. Clearly, the coefficients of the expansion in (1) are $\theta_{j\bm{\beta}}=\int_{\mathcal{X}}h_{j\bm{\beta}}(x)dQ(x)$, for all $j=1,\dots,m$.
Assuming that the true density $q$ is contained or well approximated by the alternative model in (1), a smooth test aims to assess the validity of $g_{\bm{\beta}}$ by testing the hypotheses:
$H_{0}:\theta_{1\bm{\beta}}=.=\theta_{m\bm{\beta}}=0,\text{ for some }\bm{\beta}\in\mathcal{B}\quad\quad\quad\text{versus}$ (2)
[…]
One could also rely on an exponential tilt to express the deviations of the alternative from $g_{\bm{\beta}}$ through the orthonormal set $\{h_{j\bm{\beta}}\}_{j=1}^{m}$, as in the original formulation of smooth tests proposed by Neyman, (1937). Other formulations may involve normalizing the expansion in (1) to ensure that the resulting density is non-negative (e.g., Gajek,, 1986). Here, we rely on the alternative in (1), proposed by Barton, (1953), by virtue of its simplicity. Nevertheless, the methods described in what follows can be adapted to other specifications of the alternative model for which the testing problem can be specified as in (2).
Several test statistics have been proposed in the literature to test (2). Prominent examples include the score test statistic (Neyman,, 1937; Barton,, 1953), which consists of the sum of squares of the estimated coefficients in expansion (1) and typically employed for testing simple hypotheses; the generalized score test statistic (Thomas and Pierce,, 1979; Boos,, 1992) used for testing parametric hypotheses; and the order selection test statistic (Aerts et al.,, 1999) which naturally accounts for the selection of $m$ in (1). Unfortunately, while their asymptotic null distributions are known, in practical applications, a very large sample size is typically required to reach such limits. For instance, even when $G_{\bm{\beta}}$ is the normal distribution and $m$ in (1) is fixed, the generalized score statistic requires a sample size on the order of $\sim 10^{4}$ to be reasonably approximated by its asymptotic $\chi^{2}$ distribution (Klar,, 2000). Moreover, which basis functions to be included in the expansion in (1) is typically determined by data-driven selection criteria such as the Akaike information criterion (AIC) or Bayesian information criterion (BIC) (e.g., Ledwina,, 1994; Kallenberg and Ledwina,, 1995, 1997; Inglot et al.,, 1997; Inglot and Ledwina,, 2006). In this case, the selection process introduces additional sources of variability, which affect the limiting distributions of the test statistics.
To overcome these limitations, the parametric bootstrap is often employed by practitioners to derive the null distribution of the statistics used to test (2). However, such a numerical approach may become impractical when sampling from complicated distributions and/or when the estimation of the parameters is burdensome — as it is often the case in physics and astronomy (e.g., Balázs et al.,, 2017; Cusin et al.,, 2018, 2019). Moreover, the computational cost increases substantially when testing different models since the simulations must be performed on a case-by-case basis.
This article demonstrates that most of the shortcomings arising when deriving the distributions of smooth test statistics, either asymptotically or numerically, can be overcome on the basis of recent developments in the theory of empirical processes. In particular, it is shown that the computational efficiency of the parametric bootstrap can be improved and its accuracy preserved by relying on the projected bootstrap (Cf. Section 3). Specifically, when the estimator of $\bm{\beta}$ is locally asymptotically linear, one can avoid recomputing it on each bootstrap replicate by relying on a projection of the underlying empirical process. The efficacy of this approach when deriving the distribution of classical goodness-of-fit statistics estimated via maximum likelihood has been investigated numerically by Algeri, (2022). Here, a formal proof is provided to confirm the validity of this approach in the context of smooth tests and for a more general class of estimators.
The projected bootstrap does not address scenarios in which numerical simulations are impractical due to the model’s complexity or the need to test multiple models simultaneously. Nevertheless, this limitation can be overcome by using the so-called Khmaladze-2 (K2) transform introduced by Khmaladze, (2016). In particular, we show that, in the context of smooth tests, such a transformation is especially valuable in that it enables the construction of a new set of basis functions, to be used in (1), which naturally lead to an entire new family of asymptotically distribution-free test statistics for (2). A suite of numerical studies demonstrates that the theoretical limiting null distribution of such tests well approximates the true null distribution even for samples that are only moderately large – i.e., of the order of $\sim 10^{2}$.
2 Smooth tests and the function-parametric empirical process
Let $X_{1},\ldots,X_{n}$ be independent and identically distributed (IID) random variables with CDF $Q$. Consider the Hilbert space $L^{2}(G_{\bm{\beta}})=\{h_{\bm{\beta}}:\left\langle h_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}<\infty\},$ with inner product:
$\left\langle h_{\bm{\beta}},h^{\prime}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\int_{\mathcal{X}}h_{\bm{\beta}}(x)h^{\prime}_{\bm{\beta}}(x)dG_{\bm{\beta}}(x).$
For vector-valued pairs of functions $\bm{h}_{\bm{\beta}}=[h_{1\bm{\beta}},\dots,h_{t\bm{\beta}}]^{T},\bm{h}^{{}^{\prime}}_{\bm{\beta}}=[h^{\prime}_{1\bm{\beta}},\dots,h^{\prime}_{t^{\prime}\bm{\beta}}]^{T}$ in $L^{2}(G_{\bm{\beta}})$, denote $\big{\langle}{\bm{h}_{\bm{\beta}}^{T},\bm{h}^{{}^{\prime}T}_{\bm{\beta}}\big{\rangle}}_{G_{\bm{\beta}}}$ the outer product, which results in a $t\times t^{\prime}$ matrix, with $(i,j)$-thth element given by $\int h_{i\bm{\beta}}(x)h_{j\bm{\beta}}^{{}^{\prime}}(x)dG_{\bm{\beta}}(x).$ The function-parametric empirical process $v_{G,n}$ indexed by functions $h_{\bm{\beta}}$ in $L^{2}(G_{\bm{\beta}})$ is defined as:
$v_{G,n}(h_{\bm{\beta}})=\int_{\mathcal{X}}h_{\bm{\beta}}(x)dv_{\bm{\beta},n}(x)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}h_{\bm{\beta}}(X_{i})-\sqrt{n}\big{\langle}h_{\bm{\beta}},\bm{1}\big{\rangle}_{G_{\bm{\beta}}}$
where $\bm{1}$ denotes the function identically equal to one and
$v_{\bm{\beta},n}(x)=\frac{1}{\sqrt{n}}\sum\limits_{i=1}^{n}\left[\mathds{1}_{\left\{X_{i}\leq x\right\}}-G_{\bm{\beta}}(x)\right]$
is the classical empirical process. Hence, $v_{G,n}(h_{\bm{\beta}})=v_{\bm{\beta},n}(x)$ when $h_{\bm{\beta}}(z)=\mathds{1}_{\{z\leq x\}}$.
Most statistics proposed in the literature to test (2) can be specified as functionals of the process $v_{G,n}(h_{\bm{\beta}})$ and its projections. In particular, let
$\bm{u}_{\bm{\beta}}=\nabla_{\bm{\beta}}\ln g_{\bm{\beta}}\quad\text{and}\quad\Gamma_{\bm{\beta}}=\mathbb{E}_{G_{\bm{\beta}}}\left[-\nabla_{\bm{\beta}}\bm{u}^{T}_{\bm{\beta}}\right]$
be, respectively, the score function and the Fisher information matrix of $G_{\bm{\beta}}$. Define the orthonormalized score function in $L^{2}(G_{\bm{\beta}})$ as
$\bm{b}_{\bm{\beta}}=\Gamma_{\bm{\beta}}^{-1/2}\bm{u}_{\bm{\beta}}=\left[b_{\bm{\beta}_{1}},\ldots,b_{\bm{\beta}_{p}}\right]^{T},$
where $\Gamma_{\bm{\beta}}^{-1/2}$ denotes the principal square root matrix of $\Gamma_{\bm{\beta}}^{-1}$. Denote with $\{\widetilde{h}_{j\bm{\beta}}\}_{j=1}^{m}$ the “residuals” of $\{h_{j\bm{\beta}}\}_{j=1}^{m}$ after an orthogonal projection onto $\bm{b}_{\bm{\beta}}$, i.e.,
$\widetilde{h}_{j\bm{\beta}}=h_{j\bm{\beta}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}^{T}_{\bm{\beta}},h_{j\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=h_{j\bm{\beta}}-\sum_{k=1}^{p}b_{\bm{\beta}_{k}}\left\langle h_{j\bm{\beta}},b_{\bm{\beta}_{k}}\right\rangle_{G_{\bm{\beta}}},\text{ for }j=1,\ldots,m.$ (3)
[…]
$\widehat{S}_{m,n}=\sum_{i=1}^{m}\sum_{j=1}^{m}v_{G,n}(\widetilde{h}_{i\widehat{\bm{\beta}}_{n}})\bigl{(}\widehat{\Sigma}_{m}^{-1}\bigl{)}_{ij}v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})$ (4)
where $\widehat{\Sigma}_{m}^{-1}$ is the inverse of estimated variance-covariance matrix, $\widehat{\Sigma}_{m}$, of elements
$\bigl{(}\widehat{\Sigma}_{m}\bigl{)}_{ij}=\mathbb{E}_{G_{\bm{\widehat{\beta}}_{n}}}\left[v_{G,n}(\widetilde{h}_{i\widehat{\bm{\beta}}_{n}})v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})\right]$
and $\widehat{\bm{\beta}}_{n}$ denotes a locally asymptotically linear estimator of $\bm{\beta}$, defined as in Condition (A1) in Section 3.
When the number of basis functions, $m$, to be used in (1) and (4) is determined on the basis of the data observed, the so-called order selection test statistic (cfr. Aerts et al.,, 1999) can be used to incorporate the choice of the order $m$ directly into its formulation. Specifically, let $m$ be the maximizer of the selection criterion
$\widehat{S}_{m,n}-C_{\alpha,n}m,\quad\text{ with }m=0,\dots,M_{n},$ (5)
in which $\widehat{S}_{m,n}$ is the generalized score test statistic in (4) or its non-normalized counterpart, namely,
$\widehat{S}_{m,n}=\sum_{j=1}^{m}v^{2}_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}),$ (6)
with $\widehat{S}_{0,n}=0$ in both cases. $M_{n}$ denotes the maximum number of basis functions to be potentially considered, which could either be fixed or grow to infinity as $n\to\infty.$ Selecting $m$ using the criterion in (5) corresponds to choosing the first $m$ basis functions among $\{h_{j\bm{\beta}}\}_{j=1}^{\infty}$ to be included in the expansion (1).
The underlying idea of a test based on (5) is to reject the null hypothesis in (2) if the criterion in (5) is larger than zero for some $m$ in $\mathcal{M}_{n}=\left\{1,\ldots,M_{n}\right\}$; thus, an alternative model of the form in (1) is favored over $G_{\bm{\beta}}$. Equivalently, $G_{\bm{\beta}}$ is rejected when the order selection test statistic defined as
$\widehat{T}_{n}=\max_{m\in\mathcal{M}_{n}}\left\{\frac{\widehat{S}_{m,n}}{m}\right\},$ (7)
is such that $\widehat{T}_{n}\geq C_{\alpha,n}$, for some constant $C_{\alpha,n}$ that controls the significance level, $\alpha$, of the test.
The criterion in (5) can be generalized so that basis functions indexed by any subset $B$ of $\mathcal{M}_{n}$, not necessarily ordered from $j=1,\dots,m$, are selected (Thas,, 2010, pp.103-107). In this case, the set of indices $B$ is chosen to maximize
$\widehat{S}_{B,n}-C_{n}|B|,\quad\text{ with }B\subseteq\mathcal{M}_{n},$ (8)
where $|B|$ denotes the cardinality of $B$ and $\widehat{S}_{B,n}$ is the counterpart of $\widehat{S}_{m,n}$ with the summations in (4) and (6) taken over the indexes in $B$. The corresponding subset selection test statistic $\widetilde{T}_{n}$ specifies as
$\widetilde{T}_{n}=\max_{B\subseteq\mathcal{M}_{n}:B\neq\emptyset}\left\{\frac{\widehat{S}_{B,n}}{|B|}\right\}.$ (9)
Compared to $\widehat{T}_{n}$ in (7), $\widetilde{T}_{n}$ offers greater flexibility in the selection of basis functions and thus accommodates a wider range of possible alternative models to be employed in the expansion in (1).
[…]
3 Smooth tests via projected bootstrap
The classical parametric bootstrap requires re-estimating the unknown parameter $\bm{\beta}$ on each bootstrap sample to account for the variability introduced by the estimation of the parameters. In some instances, however, a repeated estimation can make the procedure computationally intensive. The so-called “projected bootstrap” (Algeri,, 2022; Algeri and Khmaladze,, 2024) overcomes this limitation by exploiting the projection structure induced by parameter estimation – a result first established by Khmaladze, (1980) – to avoid repeating the estimation of the parameter on each bootstrap replicate. As shown in what follows, such a projection arises in the context of smooth tests rather organically.
Let the true (but unknown) parameter vector be $\bm{\beta}_{0}$, that is, under the null hypothesis $H_{0},$ the true distribution $Q$ equals $G_{\bm{\beta}_{0}}$. Let $\mathcal{N}_{\bm{\beta}_{0}}\subset\mathcal{B}$ denote the closure of a given neighborhood of $\bm{\beta}_{0}$ and denote with $\mathcal{L}\left(G_{\bm{\beta}}\right)$ the subspace of $L^{2}\left(G_{\bm{\beta}}\right)$ given by
$\mathcal{L}\left(G_{\bm{\beta}}\right)=\left\{h_{\bm{\beta}}\in L^{2}\left(G_{\bm{\beta}}\right):\left\langle h_{\bm{\beta}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=0\right\}.$
We make the following regularity assumptions:
  1. (A1)
    The estimator $\widehat{\bm{\beta}}_{n}$ has the following asymptotic representation
    $\sqrt{n}(\widehat{\bm{\beta}}_{n}-\bm{\beta}_{0})=v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}})+o_{P}(1),$
    where $\bm{\psi}_{\bm{\beta}}$ is a $p$-dimensional vector function in $\mathcal{L}\left(G_{\bm{\beta}}\right)$ and is continuously differentiable with respect to $\bm{\beta}$.
  2. (A2)
    For any $h_{\bm{\beta}}\in\mathcal{L}\left(G_{\bm{\beta}}\right)$, its gradient taken with respect to $\bm{\beta}$, denoted by $\nabla_{\bm{\beta}}h_{\bm{\beta}}$, exists and is uniformly continuous for all $x\in\mathcal{X}$ and $\bm{\beta}\in\mathcal{N}_{\bm{\beta}_{0}}$.
  3. (A3)
    For any $h_{\bm{\beta}}\in\mathcal{L}\left(G_{\bm{\beta}}\right)$ and $\bm{\beta}\in\mathcal{N}_{\bm{\beta}_{0}}$,
    $\int_{\mathcal{X}}\nabla_{\bm{\beta}}\Big{[}h_{\bm{\beta}}(x)dG_{\bm{\beta}}(x)\Big{]}=\nabla_{\bm{\beta}}\int_{\mathcal{X}}h_{\bm{\beta}}(x)dG_{\bm{\beta}}(x).$
To identify the projection structure arising from parameter estimation, we begin by demonstrating that the process $v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})$ is asymptotically equivalent to its first-order Taylor expansion around the true parameter $\bm{\beta}_{0}$.
Proposition 1.
Let $r_{n}$ be defined by the equation
$v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})=v_{G,n}(h_{\bm{\beta}_{0}})+v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}+r_{n}.$ (10)
If the assumptions (A1)-(A2) are satisfied, then $r_{n}=o_{P}(1)$.
Proof.
The result can be demonstrated by expressing $r_{n}$ as follows:
$r_{n}$ $=\sqrt{n}(\bm{\widehat{\beta}}_{n}-\bm{\beta}_{0})^{T}\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}^{{}^{\prime}}}-v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}$
$=\sqrt{n}(\bm{\widehat{\beta}}_{n}-\bm{\beta}_{0})^{T}\frac{1}{\sqrt{n}}\left[\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}^{{}^{\prime}}}-\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}\right]$
$\quad\quad+\Big{[}\sqrt{n}(\bm{\widehat{\beta}}_{n}-\bm{\beta}_{0})^{T}-v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\Big{]}\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}},$
where $\bm{\beta}$$^{{}^{\prime}}$ lies between $\bm{\beta}_{0}$ and $\widehat{\bm{\beta}}_{n}$. Under the assumptions (A1)-(A2) and by the law of large numbers, it follows that $r_{n}=o_{P}(1)$ as $n\to\infty.$
The leading terms of the right-hand side of the asymptotic representation in (10) can also be expressed in function-parametric form; as shown in Proposition 2, the functions indexing the process in (2) are projections of the functions $h_{\bm{\beta}_{0}}$.
Proposition 2.
If the assumptions (A1)-(A3) hold, then
$v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})=v_{G,n}(\Pi h_{\bm{\beta}_{0}})+o_{P}(1),$ (11)
where
$\Pi h_{\bm{\beta}}=h_{\bm{\beta}}-\bm{\psi}_{\bm{\beta}}^{T}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}$ (12)
is a projection of $h_{\bm{\beta}}$ and it is orthogonal to $\bm{u}_{\bm{\beta}}$ when $\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}=I_{p}$.
Proof.
Assumption (A3) and the law of large numbers imply
$\frac{1}{\sqrt{n}}\nabla_{\bm{\beta}}v_{G,n}(h_{\bm{\beta}})$ $=\frac{1}{\sqrt{n}}\int_{\mathcal{X}}\nabla_{\bm{\beta}}\left[h_{\bm{\beta}}(x)dv_{\bm{\beta},n}(x)\right]$ (13)
$=\frac{1}{\sqrt{n}}\int_{\mathcal{X}}[\nabla_{\bm{\beta}}h_{\bm{\beta}}(x)]dv_{\bm{\beta},n}(x)-\int_{\mathcal{X}}h_{\bm{\beta}}(x)\nabla_{\bm{\beta}}g_{\bm{\beta}}(x)dx\overset{p}{\longrightarrow}-\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}.$
Substituting (13) into (10) gives the results in (11). Additionally, when $\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}=I_{p}$,
$\Pi^{2}h_{\bm{\beta}}=h_{\bm{\beta}}-2\bm{\psi}_{\bm{\beta}}^{T}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}+\bm{\psi}_{\bm{\beta}}^{T}\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\Pi h_{\bm{\beta}},$
that is, $\Pi$ is a projection, and since
$\left\langle\bm{u}^{T}_{\bm{\beta}},\Pi h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}-\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}\left\langle\bm{u}^{T}_{\bm{\beta}},h_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=0,$
such a projection is orthogonal to $\bm{u}_{\bm{\beta}}.$
The result in equations (11)-(12) can be equivalently stated as
$v_{G,n}(h_{\widehat{\bm{\beta}}_{n}})=v_{G,n}(h_{\bm{\beta}_{0}})-v_{G,n}(\bm{\psi}_{\bm{\beta}_{0}}^{T})\left\langle\bm{u}^{T}_{\bm{\beta}_{0}},h_{\bm{\beta}_{0}}\right\rangle_{G_{\bm{\beta}_{0}}}+o_{P}(1),$
by virtue of the linearity of the projector $\Pi$. Moreover, the condition $\langle{\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\rangle}_{G_{\bm{\beta}}}=I_{p}$ is satisfied for many standard estimators, such as the maximum likelihood and the method of moments estimators. For instance, when the parameters are estimated via maximum likelihood, we have
$\bm{\psi}_{\bm{\beta}}(x)=\Gamma^{-1}_{\bm{\beta}}\bm{u}_{\bm{\beta}}(x),\quad\text{ and }\quad\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{\psi}_{\bm{\beta}}^{T}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{u}_{\bm{\beta}}^{T}\right\rangle_{G_{\bm{\beta}}}\Gamma^{-1}_{\bm{\beta}}=I_{p}.$
Furthermore, let $\mu_{k,\bm{\beta}}$ be the $k$-thth moment of $X$ under $G_{\bm{\beta}}$ and let $\bm{\varphi}_{\bm{\beta}}(x)$ be a $p$-dimensional vector-valued function with elements $x^{k}-\mu_{k,\bm{\beta}}$, for $k=1,\dots,p$. If $\widehat{\bm{\beta}}_{n}$ is the method of moments estimator, it solves the estimating equations:
$v_{G,n}(\bm{\varphi}_{\widehat{\bm{\beta}}_{n}})=0,$ (14)
and the first-order Taylor expansion of (14) around $\widehat{\bm{\beta}}_{n}=\bm{\beta}_{0}$ leads to
$\sqrt{n}(\widehat{\bm{\beta}}_{n}-\bm{\beta}_{0})=-\frac{1}{\sqrt{n}}\left[\nabla_{\bm{\beta}}v_{G,n}(\bm{\varphi}_{\bm{\beta}})\Bigl{|}_{\bm{\beta}=\bm{\beta}_{0}}\right]^{-1}v_{G,n}(\bm{\varphi}_{\bm{\beta}_{0}})+o_{P}(1),$
from which, following a similar argument as in (13), we obtain $\nabla_{\bm{\beta}}v_{G,n}(\bm{\varphi}_{\bm{\beta}})\overset{p}{\longrightarrow}-\big{\langle}{\bm{\varphi}^{T}_{\bm{\beta}},\bm{u}^{T}_{\bm{\beta}}\big{\rangle}}_{G_{\bm{\beta}}}$. Therefore,
$\bm{\psi}_{\bm{\beta}}(x)=\bigl{\langle}{\bm{\varphi}^{T}_{\bm{\beta}},\bm{u}^{T}_{\bm{\beta}}\bigl{\rangle}}_{G_{\bm{\beta}}}^{-1}\bm{\varphi}_{\bm{\beta}}(x)\quad\text{ and }\quad\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{\psi}^{T}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{u}^{T}_{\bm{\beta}},\bm{\varphi}^{T}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}\left[\left\langle\bm{\varphi}^{T}_{\bm{\beta}},\bm{u}^{T}_{\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}^{T}\right]^{-1}=I_{p}.$
In the context of smooth tests, Proposition 2 implies that the process $v_{G,n}$ indexed by the functions $\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}$ and that indexed by functions $\Pi\widetilde{h}_{j\bm{\beta}_{0}}$ are asymptotically equal. Furthermore, the functions $\widetilde{h}_{j\bm{\beta}}$ are orthogonal to $\bm{u}_{\bm{\beta}}$ by the definition in equation (3). It follows that
$\Pi\widetilde{h}_{j\bm{\beta}_{0}}=\widetilde{h}_{j\bm{\beta}_{0}}-\bm{\psi}_{\bm{\beta}_{0}}^{T}\bigl{\langle}{\bm{u}^{T}_{\bm{\beta}_{0}},\widetilde{h}_{j\bm{\beta}_{0}}\bigl{\rangle}}_{G_{\bm{\beta}_{0}}}=\widetilde{h}_{j\bm{\beta}_{0}};$
therefore, the asymptotic equality in (11) can be further simplified as in Proposition 3.
Proposition 3.
Under condition (A1), and assuming that the functions $\widetilde{h}_{j\bm{\beta}}$ satisfy the assumptions (A2) and (A3) for each $j\in\mathcal{M}_{n}$, we have
$v_{G,n}(\widetilde{h}_{j\bm{\beta}_{0}})=v_{G,n}(\Pi\widetilde{h}_{j\bm{\beta}_{0}})=v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{n}})+o_{P}(1).$ (15)
To gain some insights on the computational advantages entailed by Proposition 3, denote with $\widehat{\bm{\beta}}_{\text{obs}}$ the estimate of $\bm{\beta}$ obtained on the observed data. In the parametric bootstrap, such an estimate plays the same role as $\bm{\beta}_{0}$ in (15). Let $\widehat{\bm{\beta}}_{n}^{(b)}$ be the parameter estimate obtained on the $b$-th bootstrap sample generated from $G_{\widehat{\bm{\beta}}_{\text{obs}}}.$ Equation (15) implies that the empirical process indexed by functions $\widetilde{h}_{j\widehat{\bm{\beta}}^{(b)}_{n}}$ is asymptotically equal to the process indexed by $\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}}$. In other words, the projection structure characterizing the functions $\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}}$ makes the effect of re-estimating the parameter asymptotically negligible. Thus, for sufficiently large $n$, one only needs to evaluate the process $v_{G,n}(\widetilde{h}_{j\widehat{\bm{\beta}}_{\text{obs}}})$ over different bootstrap samples but needs not to repeat the estimation of $\widehat{\bm{\beta}}_{n}^{(b)}$ at each replicate. It follows that this projected bootstrap procedure can always make the simulation more efficient. It is particularly advantageous when dealing with computationally intensive postulated models and when parameter estimation is time-consuming. Moreover, the consistency of the projected bootstrap in recovering the true distribution of the function-parametric empirical process $v_{G,n}(\widetilde{h}_{j\bm{\beta}_{0}})$ is guaranteed under the same conditions needed for the classical parametric bootstrap (cfr. Babu and Rao,, 2004).
3.1 Empirical illustrations of smooth tests via projection
Let $G_{\bm{\beta}}$ be the asymmetric Laplace distribution with unknown asymmetry parameter $\bm{\beta}$. The corresponding PDF, $g_{\bm{\beta}}$, can be expressed as:
[…]
In what follows, the parameters $\theta$ and $\sigma$ are set equal to -10 and 2, respectively, and are assumed to be known.
A dataset of $n=100$ observations is generated from the density specified in (16), with the true value of the asymmetry parameter set to be $\bm{\beta}_{0}=0.1$. The maximum likelihood estimator $\widehat{\bm{\beta}}_{n}$ of $\bm{\beta}$ based on these observations is obtained by solving the following equation (Kotz et al.,, 2002):
$1-\frac{2\bm{\beta}^{2}}{\left(1+\bm{\beta}^{2}\right)}+\frac{\sqrt{2}}{\sigma}\left[\frac{1}{n\bm{\beta}}\sum_{j=1}^{n}\left(x_{j}-\theta\right)^{-}-\frac{\bm{\beta}}{n}\sum_{j=1}^{n}\left(x_{j}-\theta\right)^{+}\right]=0,$ (17)
where
$\left(x_{j}-\theta\right)^{-}=-\min(0,x_{j}-\theta)\quad\text{and}\quad(x_{j}-\theta)^{+}=\max(0,x_{j}-\theta).$
Given the complexity of solving (17), it is anticipated that using the projected bootstrap – in which the estimator is obtained only once rather than recalculated on each bootstrap sample – will expedite the simulation procedure.
To test the hypothesis in (2), we consider the order selection statistic in (7) and the subset selection statistic in (9). Their null distributions are simulated via three methods: the classical parametric bootstrap, the projected bootstrap, and the Monte Carlo method with data generated from the true distribution. By incorporating a simulation of Monte Carlo samples, we aim to determine whether the two bootstrap procedures considered can accurately recover the true null distribution of the test statistics for a sample size of $n=100$. All simulations are conducted using $100,000$ replicates. The orthonormal basis functions $\{h_{j\bm{\beta}}\}_{j=1}^{M_{n}}$ are chosen as compositions of the normalized shifted Legendre polynomials $h_{j}$ on $[0,1]$ with the CDF $G_{\bm{\beta}}$, i.e., $h_{j\bm{\beta}}=h_{j}\circ G_{\bm{\beta}}$. The maximum number of basis functions, $M_{n}$, is set to be 10.
Figure 1: Comparing the simulated null distributions of the order selection test statistic in (7) (Left), and subset selection test statistic in (9) (Right) via the parametric bootstrap (orange dotted lines), the projected bootstrap (grey dashed lines), and the Monte Carlo method (blue solid lines).
The results of the simulation, presented in Figure 1, demonstrate that the simulated null distributions of the order selection and subset selection test statistics are consistent across all three methods considered. This indicates that both bootstrap procedures successfully recover the true null distribution of the test statistics, even with a sample size as small as 100 observations. In terms of computational efficiency, as expected, the projected bootstrap method is significantly faster than the classical parametric bootstrap. Specifically, simulating the null distributions of the order selection and subset selection statistics using 100,000 replicates required approximately 12 minutes of CPU time using the projected bootstrap, less than a fourth of the time needed by the classical parametric bootstrap, which required approximately 51 minutes.
4 Distribution-free smooth tests via K2 transform
In general, statistics given by functionals of the process $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ are not distribution-free. Thus, while the projected bootstrap introduced in Section 3 can reduce the computational burden of simulating their null distribution, a different simulation must be implemented for each model being tested.
This section demonstrates that another route for distribution-freeness is available to practitioners. Instead of combining the values of the process $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ in a rather specific – and possibly forceful – manner to construct at most a few distribution-free statistics, the distribution-free property can be retrieved by ensuring that the empirical process $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ is itself asymptotically distribution-free under the null. Such an approach guarantees that the limiting null distribution of all its functionals is also distribution-free, thereby providing the user with an entire family of distribution-free statistics. The tool that enables such a construction is the so-called Khmaladze-2 (K2) transform [^2] The name “Khmaladze-2 transform” is used to distinguish the transformation employed in this manuscript and proposed in Khmaladze, (2016) from an earlier “Khmaladze transform” introduced by the same author in Khmaladze, (1982). introduced by Khmaladze, (2016).
In classical nonparametric goodness-of-fit testing, the probability integral transform is the transformation commonly employed to map the empirical process into the uniform empirical process, which is known to be asymptotically distributed as a standard Brownian bridge. Likewise, in the parametric setting, the K2 transform allows us to map a wide range of different projected empirical processes into standard projections that share the same limiting null distribution. When applied in the context of smooth tests, the K2 transform implicitly guides the construction of an orthonormal basis for $L^{2}(G_{\bm{\beta}})$ that guarantees the distribution-freeness of the process indexed by functions in such a basis and all its functionals.
Let $F_{\bm{\gamma}}$ be a “reference distribution” defined on the same support $\mathcal{X}$ as $G_{\bm{\beta}}$. For the moment, assume that $\bm{\gamma}$ and $\bm{\beta}$ share the same dimension; an extension to the case where $\bm{\gamma}$ has fewer parameters than $\bm{\beta}$ will be discussed in Section 4.1. The distribution $F_{\bm{\gamma}}$ constitutes the starting point of our construction and should be chosen to be simple – that is, easy to simulate from, differentiate, and evaluate. Consider the Hilbert space $L^{2}(F_{\bm{\gamma}})=\{\phi_{\bm{\gamma}}:\left\langle\phi_{\bm{\gamma}},\phi_{\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}<\infty\}$, with inner product
$\left\langle\phi_{\bm{\gamma}},\phi^{\prime}_{\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}=\int_{\mathcal{X}}\phi_{\bm{\gamma}}(x)\phi^{\prime}_{\bm{\gamma}}(x)dF_{\bm{\gamma}}(x),$
and let $\bm{a}_{\bm{\gamma}}=[a_{\gamma_{1}},\cdots,a_{{\gamma_{p}}}]^{T}$ be the orthonormalized score function of $F_{\bm{\gamma}}$. Denote with $\mathcal{L}\left(F_{\bm{\gamma}}\right)$ and $\mathcal{L}_{\perp}\left(F_{\bm{\gamma}}\right)$ the subspaces of $L^{2}\left(F_{\bm{\gamma}}\right)$ given by
$\mathcal{L}\left(F_{\bm{\gamma}}\right)=\left\{\phi_{\bm{\gamma}}\in L^{2}\left(F_{\bm{\gamma}}\right):\left\langle\phi_{\bm{\gamma}},\bm{1}\right\rangle_{F_{\bm{\gamma}}}=0\right\},\ \ \ \mathcal{L}_{\perp}\left(F_{\bm{\gamma}}\right)=\left\{\widetilde{\phi}_{\bm{\gamma}}\in\mathcal{L}\left(F_{\bm{\gamma}}\right):\bigl{\langle}{\widetilde{\phi}_{\bm{\gamma}},\bm{a}_{\bm{\gamma}}\bigl{\rangle}}_{F_{\bm{\gamma}}}=0\right\}.$
For any given orthonormal basis $\{\phi_{j\bm{\gamma}}\}_{j=1}^{\infty}$ in $\mathcal{L}(F_{\bm{\gamma}})$ the empirical process indexed by the residuals
$\widetilde{\phi}_{j\bm{\gamma}}=\phi_{j\bm{\gamma}}-\bm{a}^{T}_{\bm{\gamma}}\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}},$
is asymptotically distributed, under $F_{\bm{\gamma}},$ as a projected Brownian motion with mean and covariance given by
$\bigl{\langle}{\widetilde{\phi}_{j,\bm{\gamma}},\bm{1}\bigl{\rangle}}_{F_{\bm{\gamma}}}$ $=0\quad\text{and}$
$\bigl{\langle}{\widetilde{\phi}_{i\bm{\gamma}},\widetilde{\phi}_{j\bm{\gamma}}\bigl{\rangle}}_{F_{\bm{\gamma}}}$ $=\mathds{1}_{\{i=j\}}-\sum_{k=1}^{p}\left\langle a_{\bm{\gamma_{k}}},\phi_{i\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}\bigl{\langle}{a_{\bm{\gamma_{k}}},\phi_{j\bm{\gamma}}\bigl{\rangle}}_{F_{\bm{\gamma}}},$
respectively. Such a limiting distribution is the “standard” distribution that will be recovered through the K2 transform. In particular, since a Gaussian process is fully characterized by its mean and covariance, the K2 transform enables the construction of an orthonormal basis $\{h_{j\bm{\beta}}\}_{j=1}^{\infty}$ for $L^{2}(G_{\bm{\beta}})$, hereinafter referred to as K2 orthonormal basis, whose residuals, $\{\widetilde{h}_{j\bm{\beta}}\}_{j=1}^{\infty}$ satisfy
$\bigl{\langle}{\widetilde{h}_{j\bm{\beta}},\bm{1}\bigl{\rangle}}_{{G_{\bm{\beta}}}}=\bigl{\langle}{\widetilde{\phi}_{j\bm{\gamma}},\bm{1}\bigl{\rangle}}_{{F_{\bm{\gamma}}}}=0\quad\text{ and }\quad\bigl{\langle}{\widetilde{h}_{j\bm{\beta}},\widetilde{h}_{s\bm{\beta}}\bigl{\rangle}}_{G_{\bm{\beta}}}=\bigl{\langle}{\widetilde{\phi}_{j\bm{\gamma}},\widetilde{\phi}_{s\bm{\gamma}}\bigl{\rangle}}_{F_{\bm{\gamma}}}.$
It follows that the processes $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ and $v_{F,n}(\widetilde{\phi}_{j\bm{\gamma}})$ – as well as their functionals – have the same limiting distribution under $G_{\bm{\beta}}$ and $F_{\bm{\gamma}}$, respectively, thereby enabling the construction of a large class of asymptotically distribution-free statistics for testing (2).
For example, let $\widehat{\bm{\beta}}_{n}$ and $\widehat{\bm{\gamma}}_{n}$ denote the estimators of $\bm{\beta}$ and $\bm{\gamma}$, respectively. Consider the order-selection and subset-selection statistics defined in (7) and (9), in which $\widehat{S}_{m,n}$ chosen to be the (unnormalized) generalized score statistic in (6), with $\widetilde{h}_{j\widehat{\bm{\beta}}_{n}}$ corresponding to the residuals of the K2 orthonormal basis. Under $G_{\bm{\beta}}$, such statistics have the same limiting distribution as the statistics
$\max_{{m\in\mathcal{M}_{n}}}\sum_{j=1}^{m}\frac{v^{2}_{F,n}(\widetilde{\phi}_{j\widehat{\bm{\gamma}}_{n}})}{m}\ \ \text{ and }\ \ \max_{{B\subseteq\mathcal{M}_{n}:B\neq\emptyset}}\sum_{j\in B}\frac{v^{2}_{F,n}(\widetilde{\phi}_{j\widehat{\bm{\gamma}}_{n}})}{|B|},$ (18)
under $F_{\bm{\gamma}}$. Since the latter can be chosen arbitrarily, the limiting null distribution of the statistics in (18) can be easily simulated by means of the projected bootstrap described in Section 3. This also implies that, when testing different models, the standard limiting null distribution of the corresponding processes can be obtained using a single simulation conducted under $F_{\bm{\gamma}}$.
The construction of the K2 orthonormal basis and validation of its properties are described in detail in Section 4.1; whereas, its effectiveness in retrieving distribution-freeness in finite samples is investigated in Section 4.2 through a suite of simulation studies.
4.1 On the construction of the K2 orthonormal basis
Let $l_{\bm{\gamma},\bm{\beta}}$ be the isometry from $L^{2}(F_{\bm{\gamma}})$ to $L^{2}(G_{\bm{\beta}})$ defined as:
$l_{\bm{\gamma},\bm{\beta}}(x)=\sqrt{\frac{f_{\bm{\gamma}}(x)}{g_{\bm{\beta}}(x)}},\quad x\in\mathcal{X}.$
It can be easily verified that for all $j,$
$\left\langle l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},\bm{1}\right\rangle_{{G_{\bm{\beta}}}}\neq 0,\quad\text{and}\quad\left\langle l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},l_{\bm{\gamma},\bm{\beta}}\phi_{s\bm{\gamma}}\right\rangle_{{G_{\bm{\beta}}}}=\left\langle\phi_{j\bm{\gamma}},\phi_{s\bm{\gamma}}\right\rangle_{{F_{\bm{\gamma}}}}.$
It follows that the functions $l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$ have the same covariance as $\phi_{j\bm{\gamma}}$ under $F_{\bm{\gamma}}$, but do not share the same mean. Thus, they belong to $L^{2}(G_{\bm{\beta}})$ but not to $\mathcal{L}(G_{\bm{\beta}})$.
To rectify this, consider a linear operator, $K$, such that, when applied to $l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$, the resulting functions have mean zero under $G_{\bm{\beta}}$ but, at the same time, their covariance is preserved, that is,
$\left\langle Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},\bm{1}\right\rangle_{{G_{\bm{\beta}}}}=0,\quad\left\langle Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{s\bm{\gamma}}\right\rangle_{{G_{\bm{\beta}}}}=\left\langle\phi_{j\bm{\gamma}},\phi_{s\bm{\gamma}}\right\rangle_{{F_{\bm{\gamma}}}}.$ (19)
The second condition in (19) is satisfied for any $K$ which is unitary. For what concerns the first condition, note that
$\left\langle l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},l_{\bm{\gamma},\bm{\beta}}\right\rangle_{{G_{\bm{\beta}}}}=\left\langle\phi_{j\bm{\gamma}},\bm{1}\right\rangle_{{F_{\bm{\gamma}}}}=0;$
hence, if $Kl_{\bm{\gamma},\bm{\beta}}=\bm{1}$, then
$\left\langle Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},\bm{1}\right\rangle_{{G_{\bm{\beta}}}}=\left\langle Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},Kl_{\bm{\gamma},\bm{\beta}}\right\rangle_{{G_{\bm{\beta}}}}=\left\langle l_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},l_{\bm{\gamma},\bm{\beta}}\right\rangle_{{G_{\bm{\beta}}}}=0.$
The operator $K$ satisfying the above requirements can be the reflection
[…]
where $I$ is an identity operator. The properties of $K$ are summarized in Proposition 4 and verified in Appendix A.
Proposition 4 (Khmaladze, 2016).
The operator $K$ is unitary, and satisfies
$Kl_{\bm{\gamma},\bm{\beta}}=\bm{1},\quad K\bm{1}=l_{\bm{\gamma},\bm{\beta}},\quad\text{and if }\phi\perp span(\bm{1},l_{\bm{\gamma},\bm{\beta}}),\quad K\phi=\phi.$
Thus far, we have demonstrated that the functions $Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\in\mathcal{L}(G_{\bm{\beta}})$ have the same mean and covariance as the functions $\phi_{j\bm{\gamma}}\in\mathcal{L}(F_{\bm{\gamma}})$. Therefore, when $\bm{\gamma}$ and $\bm{\beta}$ are known, test statistics can be constructed as functionals of the empirical process $v_{G,n}(h_{j\bm{\beta}})$ and they are asymptotically distribution-free if $h_{j\bm{\beta}}=Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$.
When $\bm{\gamma}$ and $\bm{\beta}$ are unknown, an additional step is necessary to map functions $\widetilde{\phi}_{j\bm{\gamma}}\in\mathcal{L}_{\perp}(F_{\bm{\gamma}})$ into functions $\widetilde{h}_{j\bm{\beta}}$ that share the same mean and covariance. For instance, if we were to naively choose $\widetilde{h}_{j\bm{\beta}}$ to be
$\widetilde{h}_{j\bm{\beta}}=Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}^{T}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}},$
then, in general,
$\langle\widetilde{h}_{j\bm{\beta}},\widetilde{h}_{s\bm{\beta}}\rangle_{G_{\bm{\beta}}}=\left\langle\phi_{j\bm{\gamma}},\phi_{s\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}-\left\langle\bm{b}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}\left\langle\bm{b}^{T}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{s\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}\neq\langle\widetilde{\phi}_{j\bm{\gamma}},\widetilde{\phi}_{s\bm{\gamma}}\rangle_{F_{\bm{\gamma}}}.$
In the above expression, equality holds; however, if
$\big{\langle}{\bm{b}^{T}_{\bm{\beta}},Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\big{\rangle}}_{G_{\bm{\beta}}}=\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}.$
This motivates us to seek an operator, $\bm{U}_{p}$, such that, when applied to the functions $Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$, we obtain
$\left\langle\bm{b}^{T}_{\bm{\beta}},\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}},$ (21)
while ensuring their means and covariances are unchanged, that is,
$\left\langle\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=0,\quad\left\langle\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{s\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}=\left\langle\phi_{j\bm{\gamma}},\phi_{s\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}.$ (22)
Let $c_{\bm{\lambda}_{k}}$ be the kthkth component of the vector-valued function $Kl_{\bm{\gamma},\bm{\beta}}\bm{a}_{\bm{\gamma}}$, i.e.,
$c_{\bm{\lambda}_{k}}=Kl_{\bm{\gamma},\bm{\beta}}a_{\bm{\gamma}_{k}},\quad k=1,\dots,p.$
While any unitary operator $\bm{U}_{p}$ such that $\bm{U}_{p}\bm{1}=\bm{1}$ fulfills the requirements in (22), to ensure (21) holds, we need to choose $\bm{U}_{p}$ so that $\bm{U}_{p}c_{\bm{\lambda}_{k}}=b_{\bm{\beta}_{k}}$, for each $k=1,\dots,p$.
One could naively attempt to construct $\bm{U}_{p}$ as a composition of $p$ unitary operators, each mapping $c_{\bm{\lambda}_{k}}$ to $b_{\bm{\beta}_{k}}$; such an approach, however, would not lead to the desired result. To see that, consider the reflection operator on $L^{2}(G_{\bm{\beta}})$ defined as
[…]
$=I-\frac{b_{\bm{\beta}_{k}}-c_{\lambda_{k}}}{\bm{1}-\left\langle b_{\bm{\beta}_{k}},c_{\lambda_{k}}\right\rangle_{G_{\bm{\beta}}}}\left\langle b_{\bm{\beta}_{k}}-c_{\lambda_{k}},\ \cdot\ \right\rangle_{G_{\bm{\beta}}},\quad k=1,\cdots,p.$
Such an operator is self-adjoint and unitary on $L^{2}(G_{\bm{\beta}})$. It maps $c_{\bm{\lambda}_{k}}$ to $b_{\bm{\beta}_{k}}$ and $b_{\bm{\beta}_{k}}$ to $c_{\bm{\lambda}_{k}}$, while leaving all functions orthogonal to both $b_{\bm{\beta}_{k}}$ and $c_{\bm{\lambda}_{k}}$ unchanged. Let $p=2$, then
$U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{1}}=U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}b_{\bm{\beta}_{1}},$
and in general
$U_{b_{\bm{\beta}_{2}}c_{\lambda_{2}}}b_{\bm{\beta}_{1}}\neq b_{\bm{\beta}_{1}}\quad\text{unless}\quad c_{\lambda_{2}}\perp b_{\bm{\beta}_{1}}.$
To address this issue, define a set of functions $\{\widetilde{c}_{{\lambda}_{k}}\}_{k=2}^{p}$, where each $\widetilde{c}_{{\lambda}_{k}}$ is constructed to be orthogonal to $\bm{1}$ and to every $b_{\bm{\beta}_{j}}$ for which $j\leq k-1$. Specifically, we set
$\widetilde{c}_{\lambda_{k}}$ $=U_{b_{\bm{\beta}_{k-1}}\widetilde{c}_{\lambda_{k-1}}}\cdots U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{k}},\quad k=2,\ldots,p.$ (23)
A proof that each $\widetilde{c}_{\lambda_{k}}$ satisfies the required orthogonality conditions is provided in Appendix B.
Now, define the operator $\bm{U}_{p}$ as
$\bm{U}_{p}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\cdots U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}.$
Since $\bm{U}_{p}$ is a composition of unitary operators on $L^{2}\left(G_{\bm{\beta}}\right)$, it is also unitary on $L^{2}\left(G_{\bm{\beta}}\right)$. Moreover,
$\bm{U}_{p}c_{\lambda_{k}}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\cdots U_{b_{\bm{\beta}_{k}}\widetilde{c}_{\lambda_{k}}}\widetilde{c}_{\lambda_{k}}=U_{b_{\bm{\beta}_{p}}\widetilde{c}_{\lambda_{p}}}\cdots U_{b_{\bm{\beta}_{k+1}}\widetilde{c}_{\lambda_{k+1}}}b_{\bm{\beta}_{k}}=b_{\bm{\beta}_{k}},$
thus, $\bm{U}_{p}$ maps $c_{\lambda_{k}}$ into $b_{\bm{\beta}_{k}}$, for each $k=1,\dots,p$ and, since all the functions $b_{\bm{\beta}_{k}}$ and $\widetilde{c}_{\lambda_{k}}$ are orthogonal to $\bm{1}$, $\bm{U}_{p}\bm{1}=\bm{1}$. These properties are formalized in Proposition 5.
Proposition 5.
The operator $\bm{U}_{p}$ is unitary on $L^{2}\left(G_{\bm{\beta}}\right)$ and satisfies
$\bm{U}_{p}\bm{1}=\bm{1},\ \ \text{and }\ \ \bm{U}_{p}c_{\lambda_{k}}=b_{\bm{\beta}_{k}},\quad k=1,\dots,p.$
From Proposition 5, it follows that each function $\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}$ fulfills (21), and thus we obtain
$\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\widetilde{\phi}_{j\bm{\gamma}}$ $=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\left[\phi_{j\bm{\gamma}}-\bm{a}^{T}_{\bm{\gamma}}\left\langle\bm{a}^{T}_{\bm{\gamma}},\phi_{j\bm{\gamma}}\right\rangle_{F_{\bm{\gamma}}}\right]$
$=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}-\bm{b}^{T}_{\bm{\beta}}\left\langle\bm{b}^{T}_{\bm{\beta}},\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\right\rangle_{G_{\bm{\beta}}}.$
Moreover, since the operators $\bm{U}_{p},K,$ and $l_{\bm{\gamma},\bm{\beta}}$ unitary, the set of functions $\{\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}}\}_{j=1}^{\infty}$ form an orthonormal basis for $L^{2}(G_{\bm{\beta}})$, and thus we can set
$h_{j\bm{\beta}}=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\phi_{j\bm{\gamma}},\quad\text{and}\quad\widetilde{h}_{j\bm{\beta}}=\bm{U}_{p}Kl_{\bm{\gamma},\bm{\beta}}\widetilde{\phi}_{j\bm{\gamma}}.$
From the above construction, the mean and covariance of the functions $\widetilde{h}_{j\bm{\beta}}$ under $G_{\bm{\beta}}$ are the same as those of $\widetilde{\phi}_{j\bm{\gamma}}$ under $F_{\bm{\gamma}}$. Thus, the empirical processes $v_{G,n}(\widetilde{h}_{j\bm{\beta}})$ and $v_{F,n}(\widetilde{\phi}_{j\bm{\gamma}})$ have the same standard limiting distribution.
The above results can be generalized to cases where the dimension of $\bm{\gamma}$ is smaller than $p$, say $q$. In this setting, one can simply expand the orthonormal set of score functions $\{a_{\bm{\gamma}_{k}}\}_{k=1}^{q}$ to a larger orthonormal set $\{a_{\bm{\gamma}_{k}}\}_{k=1}^{p}$ in $L^{2}(F_{\bm{\gamma}})$, ensuring that all elements remain orthogonal to the constant function $\bm{1}$. This extension can be accomplished, for instance, by selecting pqp-q additional functions from another orthonormal basis in $L^{2}(F_{\bm{\gamma}})$ outside the span of $\{\bm{1},a_{\bm{\gamma}_{1}},\dots,a_{\bm{\gamma}_{q}}\}$, and applying the Gram-Schmidt orthogonalization procedure.
While at first glance the above steps may appear rather burdensome, note that all operators involved in the K2 transform are linear. Moreover, to test the hypotheses in (2), test statistics based on the K2-orthonormal basis need to be evaluated only once on the observed data for each hypothesized model $G_{\bm{\beta}}$ being tested.
4.2 Simulation Studies
Figure 2: The histogram of the simulated dataset is shown together with the densities of the true data-generating model $Q$, the reference distribution $F_{\bm{\gamma}}$, and the hypothesized distributions $G_{\bm{\beta},1},G_{\bm{\beta},2},G_{\bm{\beta},3}$. The unknown parameters $\bm{\beta}$ and $\bm{\gamma}$ are estimated via maximum likelihood.
Consider a dataset of $n=100$ observations generated from a distribution $Q$ with density
$0,10],\displaystyle q(x)=3u_{1}(x;\mu_{1},\sigma_{1})+5u_{2}(x;\mu_{2},\sigma_{2})+2u_{3}(x),\quad x\in\mathcal{X}=[-0$,0],
where $u_{1}$ and $u_{2}$ denote, respectively, the densities of truncated normal and truncated Laplace random variables; $u_{3}$ is the uniform density. The values of the location parameters are $\mu_{1}=-5$, $\mu_{2}=5$ and scale parameters are $\sigma_{1}=\sigma_{2}=3$. We aim to test the hypotheses in (2) for three different specifications of the null density, i.e.,
[…]
$g_{\bm{\beta},3}(x)$ $=\beta_{1}u_{1}(x;-4,1)+\beta_{2}u_{2}(x;4,1)+(1-\beta_{1}-\beta_{2})u_{3}(x),$
where $\bm{\beta}=(\beta_{1},\beta_{2})$ is the unknown parameter vector to be estimated.
Their corresponding CDFs are denoted by $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$, respectively. The reference distribution, $F_{\bm{\gamma}}$, is chosen to be a truncated normal distribution over $\mathcal{X}$ with unknown parameter $\bm{\gamma}$ corresponding to its mean and variance. Figure 2 shows the histogram of the dataset considered, along with the densities of $q,g_{\bm{\gamma}},g_{\bm{\beta},1},g_{\bm{\beta},2},$ and $g_{\bm{\beta},3}$ estimated via maximum likelihood.
Figure 3: The simulated null distributions of the order selection statistic (left panels) and the subset selection test statistic (right panels), using basis functions obtained by composing the normalized shifted Legendre polynomials with the null CDFs (upper panels) and the K2 transform (lower panels), under $F_{\bm{\gamma}}$, $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$.
Consider the case in which the basis functions for $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, $G_{\bm{\beta},3}$ and $F_{\bm{\gamma}}$ are constructed as compositions of the normalized shifted Legendre polynomials on [0,1] with the null CDFs. Using these basis functions, we simulated the null distributions of the order selection and subset selection test statistics in (7) and (9) with $\widehat{S}_{m,n}$ chosen to be the unnormalized generalized score statistic in (6). Specifically, the null distributions of these statistics were simulated using the projected bootstrap described in (3.1), with 100,000 replicates and a maximum of $M_{n}=6$ basis functions. As shown in the upper panel of Figure 3, the null distributions of the order selection statistic under $G_{\bm{\beta},1}$, $G_{\bm{\beta},2},$ and $G_{\bm{\beta},3}$ differs significantly from that under $F_{\bm{\gamma}}$. While the distributions of the subset selection test statistic under $F_{\bm{\gamma}}$, $G_{\bm{\beta},1}$, and $G_{\bm{\beta},3}$ are rathe similar, they differ substantially from that obtained under $G_{\bm{\beta},2}$. These discrepancies are expected given that, in general, the unnormalized generalized score statistic in (6) is not distribution-free.
$H_{0}$ $\alpha=0.001$ $\alpha=0.05$ $\alpha=0.1$
$\widehat{T}_{n}$ $\widetilde{T}_{n}$ $\widehat{T}^{K}_{n}$ $\widetilde{T}^{K}_{n}$ $\widehat{T}_{n}$ $\widetilde{T}_{n}$ $\widehat{T}^{K}_{n}$ $\widetilde{T}^{K}_{n}$ $\widehat{T}_{n}$ $\widetilde{T}_{n}$ $\widehat{T}^{K}_{n}$ $\widetilde{T}^{K}_{n}$
$F_{\bm{\gamma}}$ 0.645 0.315 0.939 0.826 0.967 0.902
$G_{\bm{\beta},1}$ 0.010 0.101 0.661 0.610 0.366 0.565 0.953 0.941 0.542 0.694 0.975 0.968
$G_{\bm{\beta},2}$ $2\times 10^{-4}$ 0.053 0.874 0.872 0.420 0.440 0.991 0.989 0.570 0.576 0.996 0.995
$G_{\bm{\beta},3}$ 0.005 0.026 0.037 0.026 0.159 0.266 0.349 0.302 0.277 0.383 0.469 0.427
Table 1: Comparing the power of the classical order selection and subset selection test statistics with their K2-transformed counterparts. The significance levels are $0.001,0.05,$ or $0.1$.
The same experiment was repeated, considering the K2 transformed basis functions for $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, $G_{\bm{\beta},3}$ constructed as described in Section 4.1. As shown in the lower panels of Figure 3, for both statistics considered, the simulated null distributions under $G_{\bm{\beta},1}$, $G_{\bm{\beta},2}$, and $G_{\bm{\beta},3}$ are all indistinguishable from those under $F_{\bm{\gamma}}$. It follows that, for all models and statistics considered in this example, a sample size of 100 observations is sufficient to retrieve the distribution-free property even when two parameters are estimated.
Table 1 compares the power of test statistics constructed using the K2 basis functions with that obtained when considering compositions of normalized shifted Legendre polynomials with the null CDF. For all models under study, the K2-based test statistics exhibit higher power. However, this result should not be assumed to be true in general: the K2 transform leads to the construction of a new family of test statistics that can outperform the classical ones in certain settings but are not guaranteed to always lead to higher power.
5 Case study: analyzing an X-ray spectrum from RT Cru
In X-ray astronomy, spectral analysis is essential for understanding the fundamental properties of stars, galaxies, and other celestial objects. In particular, the presence of spectral lines in X-ray spectra provides valuable insights into an object’s chemical composition, distance from Earth, temperature, motion, surrounding environments, and other important attributes.
[…]
$g_{\bm{\beta}}(x)=(1-\beta_{1}-\beta_{2}-\beta_{3})b(x)+\sum\limits_{r=1}^{3}\beta_{r}s_{r}(x),\quad x\in\mathcal{X},$ (24)
where $\mathcal{X}=[1.65,2.05]$; $b(x)$ is a uniform background density on $\mathcal{X}$; and the functions $s_{r}(x)$ model each of the the three expected spectral lines. They consist of a convolution of a normal density with a Moffat function [^4] The Moffat function is equivalent to the density of a location-scale Student’s $t$-distribution with a location of 0, a scale of 0.025, and 4 degrees of freedom. (Moffat,, 1969) and specify as
$s_{r}\left(x\right)\propto\int_{-\infty}^{+\infty}\frac{\exp\left\{-\frac{1}{2}\left(\frac{w_{0}-\mu_{r}}{\sigma_{r}}\right)^{2}\right\}}{\left[1+\left(\frac{x-w_{0}}{0.05}\right)^{2}\right]^{2.5}}dw_{0},\quad x\in\mathcal{X},$ (25)
where $\mu_{r},\sigma_{r}$ are known parameters, with $\mu_{1}=1.78499,\mu_{2}=1.85247$, and $\mu_{3}=1.94365$, and $\sigma_{1}=\sigma_{2}=\sigma_{3}=0.0025$. The unknown parameters $\beta_{r}$ represent the relative intensities of the spectral lines and are estimated via maximum likelihood.
Figure 4: Left: $Q$ plots of the simulated order selection test statistics in (7) under $G_{\bm{\beta}}$ and in (18) under $F_{\bm{\gamma}}$. Right: $Q$ plots of the simulated subset selection test statistics in (9) under $G_{\bm{\beta}}$ and in (18) under $F_{\bm{\gamma}}$.
We use the approach described in Section 4 to test the validity of (24). The reference distribution $F_{\bm{\gamma}}$ considered is a simplified version of (24). In particular, it consists of a convex combination of the uniform distribution and truncated normal distributions over $\mathcal{X}$. Its density is:
[…]
where $p_{r}$ is the density of the truncated normal over $\mathcal{X}$, with $\mu_{r}$ being the known positions of the spectral lines and given after (25). The unknown parameter $\bm{\gamma}$ is estimated via the maximum likelihood. The basis functions for $F_{\bm{\gamma}}$ are chosen to be compositions of normalized shifted Legendre polynomials on $[0,1]$ with $F_{\bm{\gamma}}$. These functions are employed as a starting point to construct the K2 transformed basis functions $h_{j\bm{\beta}}$ for $G_{\bm{\beta}}$, whose residuals are subsequently used when calculating the order selection and subset selection statistics in (7) and (9), respectively, with $\widehat{S}_{m,n}$ as in (6). Their limiting null distributions are obtained by simulating 100,000 realizations of the statistics (18) under $F_{\bm{\gamma}}$ through the projected bootstrap (Cf. Section 3). The maximum number of basis functions used is $M_{n}=6$. To assess the accuracy of the approximation, we also simulated the null distributions of the statistics in (7) and (9) constructed using the K2 basis functions directly from $G_{\bm{\beta}}$. As shown by the QQ-plots in Figure 4, the resulting null distributions of (7) and (9) under $G_{\bm{\beta}}$ are indistinguishable from those of (18) under $F_{\bm{\gamma}}$.
Finally, the p-values from the order selection and subset selection test statistics are 0.463 and 0.454, respectively, indicating that the hypothesized model in (24) fits the observed spectrum well.
[…]
6 Summary and discussion
This article introduces a novel framework for constructing asymptotically distribution-free smooth tests that do not rely on the usual $\chi^{2}$ approximation. It is shown that, even when the parameters are estimated, the asymptotic null distribution provides a reasonable approximation when the sample size is only moderately large.
The asymptotic distribution-free property is achieved by relying on the K2 transform. The latter consists of a change of variable in functional space, which enables the construction of an empirical process with a standard asymptotic null distribution. In the context of smooth tests, such a transformation is especially valuable in that it yields a new family of orthonormal bases in $L^{2}(G_{\beta})$ such that, even when the parameters are estimated, test statistics defined by functions in these bases are asymptotically distribution-free.
The projected bootstrap is also discussed as a computationally efficient alternative to the classical parametric bootstrap. In particular, the projection structure induced by parameter estimation allows us to simulate the null distribution of the test statistics of interest without re-estimating the model parameters at each bootstrap replicate. Simulation experiments show that the computational gain attained by projected bootstrap can be substantial compared to the parametric bootstrap, especially when the estimation of the parameters is CPU-intensive.
While the present manuscript focuses on the univariate setting, the proposed framework can be easily adapted to test multivariate parametric models. In particular, when $G_{\bm{\beta}}$ is $D$-dimensional distribution, the reference distribution, $F_{\bm{\gamma}}$ could be chosen to be a product of $D$ univariate distributions $F_{\bm{\gamma},d}$, $d=1,\dots,D$. The $K2$ orthonormal basis in $L^{2}(G_{\bm{\beta}})$ can then be constructed by applying the K2 transform to a tensor product of bases in $L^{2}(F_{\gamma,d})$ (e.g., Algeri, (2021)).
Extensions to smooth tests for regression (Rayner et al.,, 2022) are also possible. In this case, instead of relying on the classical empirical process for i.i.d. data to express the statistics of interest as functionals from it (see Section 2), the random measure to be used is the weighted empirical process proposed in Khmaladze, (2017).
7 Acknowledgments
The authors thank Charlie Geyer for the valuable feedback.
The work of XZ and SA was in part supported by the Research and Innovation Office at the University of Minnesota. SA was partially supported by NSF grant DMS-2152746. XZ was partially supported by the University of Minnesota Data Science Initiative.
8 Code Availability
The R code used to conduct the simulations and analyses in Sections 3.1, 4.2, and 5 is available at https://github.com/xiangyu2022/DisfreeSmoothTests.
Appendix
A Proof of Proposition 4
Proof.
To show that $K$ is a unitary operator, we need to demonstrate that it is surjective and preserves the inner product. It is surjective because for any function in $L^{2}(G_{\bm{\beta}})$, if $\phi\perp span(\bm{1},l_{\bm{\gamma},\bm{\beta}})$,
$K\phi$ $=\phi-\frac{l_{\bm{\gamma},\bm{\beta}}-\bm{1}}{\bm{1}-\langle l_{\bm{\gamma},\bm{\beta}},\bm{1}\rangle_{G_{\bm{\beta}}}}\langle l_{\bm{\gamma},\bm{\beta}}-\bm{1},\phi\rangle_{G_{\bm{\beta}}}=\phi.$
Otherwise if $\phi=c_{1}\bm{1}+c_{2}l_{\bm{\gamma},\bm{\beta}},$ we have
$K(c_{1}l_{\bm{\gamma},\bm{\beta}}+c_{2}\bm{1})$ $=c_{1}l_{\bm{\gamma},\bm{\beta}}+c_{2}\bm{1}-\frac{l_{\bm{\gamma},\bm{\beta}}-\bm{1}}{\bm{1}-\langle l_{\bm{\gamma},\bm{\beta}},\bm{1}\rangle_{F_{\bm{\gamma}}}}\langle l_{\bm{\gamma},\bm{\beta}}-\bm{1},c_{1}l_{\bm{\gamma},\bm{\beta}}+c_{2}\bm{1}\rangle_{F_{\bm{\gamma}}}$ (26)
$=c_{1}l_{\bm{\gamma},\bm{\beta}}+c_{2}\bm{1}-c_{1}(l_{\bm{\gamma},\bm{\beta}}-\bm{1})+c_{2}(l_{\bm{\gamma},\bm{\beta}}-\bm{1})$
$=c_{1}\bm{1}+c_{2}l_{\bm{\gamma},\bm{\beta}}.$
Notice here, by letting $c_{1}=1,c_{2}=0$ or $c_{1}=0,c_{2}=1$ in equation (26), we obtain
$Kl_{\bm{\gamma},\bm{\beta}}=\bm{1},\quad K\bm{1}=l_{\bm{\gamma},\bm{\beta}}.$
The operator $K$ preserves the inner product because for any functions $\phi_{1},\phi_{2}\in L^{2}(G_{\bm{\beta}})$
$\left\langle K\phi_{1},K\phi_{2}\right\rangle_{G_{\bm{\beta}}}$
$=\left\langle\phi_{1},\phi_{2}\right\rangle_{G_{\bm{\beta}}}-\frac{2\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\phi_{1}\right\rangle_{G_{\bm{\beta}}}\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\phi_{2}\right\rangle_{G_{\bm{\beta}}}}{1-\left\langle l_{\bm{\gamma},\bm{\beta}},1\right\rangle_{G_{\bm{\beta}}}}$
$\hskip 62.59596pt+\frac{\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\phi_{1}\right\rangle_{G_{\bm{\beta}}}\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\phi_{2}\right\rangle_{G_{\bm{\beta}}}\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\bm{1}-l_{\bm{\gamma},\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}}{(1-\left\langle l_{\bm{\gamma},\bm{\beta}},1\right\rangle_{G_{\bm{\beta}}})^{2}}$
$=\left\langle\phi_{1},\phi_{2}\right\rangle_{G_{\bm{\beta}}}+\frac{\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\phi_{1}\right\rangle_{G_{\bm{\beta}}}\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\phi_{2}\right\rangle_{G_{\bm{\beta}}}}{(1-\left\langle l_{\bm{\gamma},\bm{\beta}},1\right\rangle_{G_{\bm{\beta}}})^{2}}\left(-2+2\left\langle l_{\bm{\gamma},\bm{\beta}},1\right\rangle_{G_{\bm{\beta}}}+\left\langle\bm{1}-l_{\bm{\gamma},\bm{\beta}},\bm{1}-l_{\bm{\gamma},\bm{\beta}}\right\rangle_{G_{\bm{\beta}}}\right)$
$=\left\langle\phi_{1},\phi_{2}\right\rangle_{G_{\bm{\beta}}}.$
$K$ is self-adjoint because
$\left\langle K\phi_{1},\phi_{2}\right\rangle_{G_{\bm{\beta}}}$ $=\left\langle\phi_{1}-\frac{l_{\bm{\gamma},\bm{\beta}}-\bm{1}}{\bm{1}-\langle l_{\bm{\gamma},\bm{\beta}},\bm{1}\rangle_{G_{\bm{\beta}}}}\langle l_{\bm{\gamma},\bm{\beta}}-\bm{1},\phi_{1}\rangle_{G_{\bm{\beta}}},\phi_{2}\right\rangle_{G_{\bm{\beta}}}$
$=\left\langle\phi_{1},\phi_{2}\right\rangle_{G_{\bm{\beta}}}-\frac{\left\langle l_{\bm{\gamma},\bm{\beta}}-\bm{1},\phi_{1}\right\rangle_{G_{\bm{\beta}}}\left\langle l_{\bm{\gamma},\bm{\beta}}-\bm{1},\phi_{2}\right\rangle_{G_{\bm{\beta}}}}{\bm{1}-\langle l_{\bm{\gamma},\bm{\beta}},\bm{1}\rangle_{G_{\bm{\beta}}}}$
$=\left\langle\phi_{1},\phi_{2}-\frac{l_{\bm{\gamma},\bm{\beta}}-\bm{1}}{\bm{1}-\langle l_{\bm{\gamma},\bm{\beta}},\bm{1}\rangle_{G_{\bm{\beta}}}}\langle l_{\bm{\gamma},\bm{\beta}}-\bm{1},\phi_{2}\rangle_{F_{\bm{\gamma}}}\right\rangle_{G_{\bm{\beta}}}$
$=\left\langle\phi_{1},K\phi_{2}\right\rangle_{G_{\bm{\beta}}}.$
The unitary and self-adjoint properties of $K$ imply $K^{2}=I.$
B Required Orthogonality Conditions of $\widetilde{c}_{\lambda_{k}}$
Proof.
It can be verified that
$\left\langle\widetilde{c}_{\lambda_{2}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{2}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{2}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{2}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=0,$
$\left\langle\widetilde{c}_{\lambda_{2}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{2}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{2}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{2}},c_{\lambda_{1}}\right\rangle_{G_{\bm{\beta}}}=0;$
and
$\left\langle\widetilde{c}_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}\bm{1}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},\bm{1}\right\rangle_{G_{\bm{\beta}}}=0,$
$\left\langle\widetilde{c}_{\lambda_{3}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}b_{\bm{\beta}_{1}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},c_{\lambda_{1}}\right\rangle_{G_{\bm{\beta}}}=0,$
$\left\langle\widetilde{c}_{\lambda_{3}},b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}=\left\langle U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{3}},b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}U_{b_{\bm{\beta}_{2}}\widetilde{c}_{\lambda_{2}}}b_{\bm{\beta}_{2}}\right\rangle_{G_{\bm{\beta}}}=\left\langle c_{\lambda_{3}},U^{2}_{b_{\bm{\beta}_{1}}c_{\lambda_{1}}}c_{\lambda_{1}}\right\rangle_{G_{\bm{\beta}}}=0;$
and this can proceed up to $\widetilde{c}_{\lambda_{p}}$ by induction.